A stunt driver wants to make his car jump over 10 cars parked side by side below a horizontal
ramp. With what minimum speed must he drive off the ramp? The vertical height of
the ramp is 2.0 m, and the horizontal distance he must clear is 24 m.
1
Expert's answer
2019-09-19T09:36:42-0400

The horizontal distance traveled by car


"L=v_0\sqrt{\frac{2H}{g}}"

So, the minimum speed


"v_0=L\sqrt{\frac{g}{2H}}=24\:\rm m\times \sqrt{\frac{9.8\:\rm m/s^2}{2\times 2.0\:\rm m}}=37.6\:\rm m/s"